Self - Assessment

A. List the members of the following sets

1. {x| x is real numbers and x2 = 1}

2. {x| x is an integer and -4 < x ≤ 3}

B. Use set builder notation to give description of each of these sets.

1. {a, e,i ,o, u}

2. {=2, -1, 0, 1, 2}

C. Let A= (a, b, c), B = (x, y) and C = (0, 1)

Find:

1. A U C

2. C x B

3. B – A

4. (A ∩ C) U B

D. Find these terms of the sequence (An}, where An = 2(3)n + 5

1. A0

2. A5

3. A3

4. 8th term

5. 2nd term

6. Sum of the sequence

E. Given the following set:

2. X = {-1, 0, 1, 2, 3, 4, 5} defined by the rule (x, y) ∈R if x ≤ y

F. List the elements of R

G. Find the domain of R

H. Find the range of R

I. Draw the digraph

J. Properties of the Relation


1
Expert's answer
2021-03-18T15:48:58-0400

Solution:

(A):

1: "\{ 1,-1\}"

2: "\{ -3,-2,-1,0,1,3\}"

(B):

1: {x| x is a vowel }

2: {x| x is an integer and "-2\le x\le2" }

(C):

1: "A\cup C=\{ a,b,c,0,1\}"

2: "C\times B=\{ (0,x),(0,y),(1,x),(1,y)\}"

3: "B-A=\{ x,y\}"

4: "(A\cap C)\cup B=\{\phi\}\cup \{x,y\}=\{x,y\}"

(D):

"A_n=2(3)^n+5"

(1): Put n = 0

"A_0=2(3)^0+5=2(1)+5=7"

(2): Put n = 5

"A_5=2(3)^5+5=2(243)+5=491"

(3): Put n = 3

"A_3=2(3)^3+5=2(27)+5=59"

(4): For 8th term, put n = 8

"A_8=2(3)^8+5=2(6561)+5=13127"

(5): For 2nd term, put n = 2

"A_2=2(3)^2+5=2(9)+5=23"

(6): "S_n=\sum [2(3)^n+5]=2\sum 3^n+5\sum 1"

"=2(3^0+3^1+...+3^n)+5n
\\=2[\dfrac{1(3^{n-1}-1)}{3-1}]+5n \ [\text{Using GP}] 
\\=3^{n-1}+5n-1"

(E):

(F): R = {(-1,-1),(-1,0),(-1,1),(-1,2),(-1,3),(-1,4),(-1,5),(0,0),(0,1),(0,2),(0,3),(0,4),(0,5),(1,1),(1,2),(1,3),(1,4),(1,5),(2,2),(2,3),(2,4),(2,5),(3,3),(3,4),(3,5),(4,4),(4,5),(5,5)}

(G): Domain of R = {-1, 0, 1, 2, 3, 4, 5}

(H): Range of R = {-1, 0, 1, 2, 3, 4, 5}

(I): Digraph:



(J): This relation is reflexive and transitive but not symmetric as-

Reflexive: "(x\le x)", this is true

Transitive: "\\(x\le y)\ \& (y\le z)\Rightarrow (x\le z)", this is true

Symmetric: "(x\le y)\Rightarrow (y\le x)", this is not true.